### v2

# Probabilistic Construction of Kakeya-Type Sets in $\mathbb{R}^{2}$ associated to separated sets of directions

Abstract.
We provide a condition on a set of directions $\Omega\subset\mathbb{S}^{1}$ ensuring that the associated directional maximal operator $M_{\Omega}$ is unbounded on $L^{p}(\mathbb{R}^{2})$ for every $1\leq p<\infty$. The techniques of proof extend ideas of Bateman and Katz involving probabilistic construction of Kakeya-type sets using sticky maps and Bernoulli percolation.
Key words and phrases:
maximal functions, differentiation basis
2020 Mathematics Subject Classification:
Primary 42B25
P. H. is partially supported by a grant from the Simons Foundation (#521719 to Paul Hagelstein).
1. Introduction
This paper addresses problems associated to the $L^{p}(\mathbb{R}^{2})$ boundedness of directional maximal operators acting on measurable functions on $\mathbb{R}^{2}$. In particular we provide a condition on a set of directions so that the associated maximal operator is unbounded on $L^{p}(\mathbb{R}^{2})$ for every $1\leq p<\infty$. Our research extends the classical work of Nikodym [10] and Busemann and Feller [3], who constructed Kakeya-type sets that may be used to provide examples indicating that the directional maximal operator associated to the set of all directions in $\mathbb{S}^{1}$ is unbounded on $L^{p}(\mathbb{R}^{2})$ for every $1\leq p<\infty$. It more closely relates, however, to the more recent work of Bateman and Katz [2] and Bateman [1] that indicates how probabilistic techniques may be used to show that certain directional maximal operators are unbounded on $L^{p}(\mathbb{R}^{2})$ for every $1\leq p<\infty$. A particularly noteworthy result in [2] in this regard due to Bateman and Katz is that if $\Omega$ is the Cantor ternary set in $[0,1]$, then the associated directional maximal operator $M_{\Omega}$ acting on measurable functions in $\mathbb{R}^{2}$ is unbounded on $L^{p}(\mathbb{R}^{2})$ for all $1\leq p<\infty$. The goal of this paper is to show that the primary ideas of the paper of Bateman and Katz may, with appropriate modifications, yield similar results for sets that are not lacunary of finite order but satisfy a certain “separation” condition.
The paper [1] contains a theorem asserting that, if $\Omega$ is a subset of $\mathbb{S}^{1}$ that is not the union of finitely many sets of finite lacunary order, then the associated maximal operator $M_{\Omega}$ is not bounded on $L^{p}(\mathbb{R}^{2})$ for any $1\leq p<\infty$. We have recently uncovered a subtle quantitative error in the proof of this theorem that is discussed in Section 4 of this paper. At the present time, to the best of our knowledge, the correctness of the statement of this theorem is unknown. That being said, extension and modification of techniques in [1] do enable us to assert for a wide class of sets of directions $\Omega\subset\mathbb{S}^{1}$ that the associated directional maximal operators $M_{\Omega}$ are unbounded on $L^{p}(\mathbb{R}^{2})$ for every $1\leq p<\infty$.
In our paper, we will associate to a given set of directions $\Omega\subset\mathbb{S}^{1}$ a lacunary value $\lambda(\Omega)$. The definition of the lacunary value $\lambda(\Omega)$ will be very much in the spirit of Bateman’s paper [1]. In addition to the lacunary value $\lambda(\Omega)$ associated to a given set of directions $\Omega$, we will introduce the notion that a set of directions is $\eta$-separated. Loosely speaking, we would say that the ternary Cantor set is $\frac{1}{3}$-separated as the distance between the intervals $[0,\frac{1}{3}]$ and $[\frac{2}{3},1]$ is $\frac{1}{3}$ the length of the ambient interval $[0,1]$, with a similar relation holding for subsequent intervals in the natural construction of the ternary Cantor set. This positive ratio is crucial in the Bateman and Katz proof that the directional maximal operator associated to a Cantor set is unbounded on $L^{p}(\mathbb{R}^{2})$ for all $1\leq p<\infty$. However, as we shall see, this positive ratio does not exist for general sets of infinite lacunary value, prohibiting the type of Bernoulli $(\frac{1}{2})$ percolation argument used by Bateman and Katz in [2] to also be used in the same manner to show that if $\Omega$ is a set of directions in $\mathbb{S}^{1}$ with $\lambda(\Omega)=\infty$, then $M_{\Omega}$ is necessarily unbounded on $L^{p}(\mathbb{R}^{2})$ for every $1\leq p<\infty$. The main result in our paper is that, if $\Omega\subset\mathbb{S}^{1}$ is such that, for some $\eta>0$, $\Omega$ contains $\eta$-separated subsets $\Omega_{N}$ with $\lambda(\Omega_{N})=N$ for every natural number $N$, then $M_{\Omega}$ is unbounded on $L^{p}(\mathbb{R}^{2})$ for every $1\leq p<\infty$.
The organization of the paper is as follows. In the second section we will define certain terminology used in the paper, indicating what we mean by the lacunary value $\lambda(\Omega)$ of a set $\Omega\subset\mathbb{S}^{1}$ and the directional maximal operator $M_{\Omega}$ associated to $\Omega$. We will also define the $\eta$-separation condition. In this section we will state the main theorem of the paper as well as provide an overview of the structure of the main theorem, indicating, motivated by Bateman’s paper, that there exist positive constants $c_{\eta}$ and $C_{\eta,N}$ so that $\lim_{N\rightarrow\infty}C_{\eta,N}=\infty$ and so that if $\Omega$ contains an $\eta$-separated subset of lacunary value $N$, then there exist sets $K_{1}$ and $K_{2}$ in $\mathbb{R}^{2}$ constructed probabilistically such that $|K_{1}|\geq C_{\eta,N}|K_{2}|$ and such that $M_{\Omega}\chi_{K_{2}}>c_{\eta}$ on $K_{1}$. In this section we will recall lower estimates on the measures of all $K_{1}$-type sets as provided by Bateman. Section 3 will be devoted to the probabilistic construction of a $K_{2}$-type set whose measure satisfies a desired upper estimate. In Section 4 we will provide, given $N$, an example of a set $\Omega\subset\mathbb{S}^{1}$ that is $N$-lacunary but such that, letting $\mathscr{T}_{\Omega}$ be the subset associated to $\Omega$ of the binary tree and defining for each sticky map $\sigma:\mathscr{B}^{h(\mathscr{T}_{\Omega})}\rightarrow\mathscr{T}_{\Omega}$ the associated set $K_{\sigma}$ as in [1], we have $\sup_{(x,y)\in\mathbb{R}^{2}\atop{x\geq 1}}Pr((x,y)\in K_{\sigma})=1$, where the probability is taken over all such sticky maps. This provides a counterexample to a step in the proof of Claim 7(B) of [1] which asserted that for all $x\geq 1$ one has $Pr((x,y)\in K_{\sigma})\lesssim\frac{1}{N}$. In Section 5 we will provide an example of a set $\Omega\subset\mathbb{S}^{1}$ that, although having infinite lacunary value, for no $\eta>0$ contains an $\eta$-separated set of lacunary value $N$ for every finite value of $N$. This example, however, does not provide a counterexample to Claim 7(B) itself. Additionally we will make concluding remarks and make suggestions for further research in this area.
We wish to express our gratitude to the referees for their comments and suggestions regarding this paper.
2. Terminology and Statement of Main Theorem
Let $\Omega$ be a nonempty subset of $\mathbb{S}^{1}$. We may associate to $\Omega$ the directional maximal operator $M_{\Omega}$ acting on measurable functions on $\mathbb{R}^{2}$ by
$M_{\Omega}f(x):=\sup_{x\in R}\frac{1}{|R|}\int_{R}|f|\;\textup{,}$
where the supremum is taken over the set of all open rectangles in $\mathbb{R}^{2}$ containing $x$ with an edge of longest length being oriented in one of the points (directions) of $\Omega$.
For the remainder of the paper we will assume, without loss of generality, that $\Omega\subset\mathbb{S}^{1}$ is such that, for every $\omega\in\Omega$, the line $\ell\subseteq\mathbb{R}^{2}$ passing through the origin and $\omega$ intersects the line segment
$\{(1,u):0\leq u\leq 1\}\;\textup{}$
at a point $\omega_{Q}$. For convenience, we will identify $\Omega$ with the set
$Q_{\Omega}:=\{u:(1,u)=\omega_{Q}\textup{ for some }\omega\in\Omega\}\;$
or more simply identify $\Omega$ with a subset of $[0,1]$. (Equivalently, we can assume $\Omega$ lies in the first octant of the plane and we identify $\Omega$ with the tangents of the associated angles.)
We now indicate how we will denote dyadic subintervals of $[0,1]$. Let $Q_{0}$ denote the interval $[0,1]$. Let $Q_{00},Q_{01}$ denote the two closed a.e. disjoint dyadic subintervals of $Q_{0}$ whose union forms $Q_{0}$, where all the values in $Q_{00}$ are less than or equal to any value in $Q_{01}$. Continuing recursively, given $j_{i}\in\{0,1\}$, for $1\leq i\leq k$, we let $Q_{0j_{1}\ldots j_{k}0},Q_{0j_{1}\ldots j_{k}1}$ denote the two nonoverlapping dyadic subintervals of $Q_{0j_{1}\ldots j_{k}}$ whose union forms $Q_{0j_{1}\ldots j_{k}}$, where all the values in $Q_{0j_{1}\ldots j_{k}0}$ are less than or equal to any value in $Q_{0j_{1}\ldots j_{k}1}$. If $u$ is the binary string $0j_{1}j_{2}\ldots j_{k}$, we may abbreviate the interval $Q_{0j_{1}j_{2}\ldots j_{k}}$ by $Q_{u}$. When convenient, we will also let $u=0j_{1}j_{2}\ldots j_{k}$ denote the interval $[\sum_{i=1}^{k}2^{-i}j_{i},\sum_{i=1}^{k}2^{-i}j_{i}+2^{-k}].$
We define the binary tree $\mathscr{B}$ to be the graph whose vertex set consists of 00 and all finite strings of the form $0a_{1}a_{2}\ldots a_{k}$ where each $a_{i}\in\{0,1\}$, and whose edge set is the collection of unordered pairs of vertices of the form $(0,0a_{1})$ or $(0a_{1}\ldots a_{k-1},0a_{1}\ldots a_{k-1}a_{k})$.
Given $\Omega\subset\mathbb{S}^{1}$, we define $\mathscr{T}_{\Omega}$ to be the smallest subtree of $\mathscr{B}$ containing 0 and all of the vertices of the form $0a_{1}a_{2}\ldots a_{k}$ such that $Q_{0a_{1}a_{2}\ldots a_{k}}\cap Q_{\Omega}\neq\emptyset$.
Let $\mathscr{T}$ be a subtree of $\mathscr{B}$. Any vertex $v\in\mathscr{T}$ of the form $v=0a_{1}\ldots a_{k}$ is said to be of height $k$, and we may write $h(v)=k$. $0\in\mathscr{T}$ is considered to be of height 00. The height of a nonempty tree is the supremum of the heights of its vertices. If $u\textup{,}v\in\mathscr{T}$, an edge in $\mathscr{B}$ exists connecting $u$ and $v$, and $h(v)=1+h(u)$, $u$ is considered to be a parent of $v$ and $v$ is considered to be a child of $u$. If $u_{j}$ is a parent of $u_{j+1}$ for $j=0,\ldots,k-1$, then $u_{j}$ is considered to be an ancestor of $u_{k}$ and $u_{k}$ is considered to be a descendant of $u_{j}$. If the vertex $u\in\mathscr{T}$ has two children in $\mathscr{T}$, then the vertex $u$ is considered to split in $\mathscr{T}$, and we may also refer to $u$ as a splitting vertex.
If $\mathscr{T}$ is a subtree of $\mathscr{B}$ and $N$ is a natural number, we define $\mathscr{T}^{N}$ to be the truncation of $\mathscr{T}$ to all of its vertices of height less than or equal to $N$.
A ray $R$ in $\mathscr{T}$ is a (possibly infinite) maximal ordered collection of vertices $v_{1},v_{2},v_{3},\ldots$ in $\mathscr{T}$ such that $h(v_{j+1})=1+h(v_{j})$. It is maximal in the sense that the ray does not terminate at a vertex $v\in\mathscr{T}$ if $v$ has any descendants in $\mathscr{T}$. If $v\in\mathscr{T}$, the set of rays starting at $v$ of the form $v,v_{2},v_{3},\ldots$ is labeled by $\mathfrak{R}_{\mathscr{T}}(v)$.
Given a tree $\mathscr{T}$ and a ray $R$ in $\mathscr{T}$, we define the splitting number $\textup{split}(R)$ of $R$ to be the number (possibly infinite) of vertices that split in $\mathscr{T}$ that lie on $R$. The splitting number of a vertex $v$ with respect to a tree $\mathscr{S}$ rooted at $v$ is defined by
$\textup{split}_{\mathscr{S}}(v):=\min_{R\in\mathfrak{R}_{\mathscr{S}}(v)}\textup{split}(R)\;.$
The splitting number of a vertex $v$ in a tree $\mathscr{T}$ is defined by
$\textup{split}(v):=\sup_{\mathscr{S}\subset\mathscr{T}}\textup{split}_{\mathscr{S}}(v)\;,$
where the supremum is over all subtrees $\mathscr{S}$ of $\mathscr{T}$ all of whose vertices are of height at least that of $v$. We define
$\textup{split}(\mathscr{T}):=\sup_{v\in\mathscr{T}}\textup{split}(v)\;.$
Given $\Omega\subset\mathbb{S}^{1}$, we define the lacunary value $\lambda(\Omega)$ by
$\lambda(\Omega):=\textup{split}(\mathscr{T}_{\Omega})\;.$
Although this terminology is motivated by that of Sjögren and Sjölin [11] and Bateman [1], a few words of caution are in order here. To begin with, the lacunary value does not agree with what is typically considered the lacunary order of a set. As an example, if $\Omega=\{1/2\}$, then $\lambda(\Omega)=1$ since $1/2$ has a binary representation of both 100000… and 0111…. Similarly, multiple binary representations of numbers of the form $1/2^{j}$ lead us to have that the lacunary value of the set $\{1/2,1/4,1/8,\ldots\}$ is 2 although this set is typically considered to have lacunary order 1. It is for this reason that we refer to a lacunary value of a set as opposed to a lacunary order. We suppose we could get around this issue by associating to any point in $\Omega$ a single ray, say by choosing a ray that was minimal with respect to a type of dictionary order, but this would create a certain degree of artificiality that we wish to avoid. At any rate, the lacunary value $\lambda(\Omega)$ that we define agrees with the splitting number $\textup{split}(\mathscr{T}_{\Omega})$ defined by Bateman, so our definition would seem to be reasonable.
Again following terminology in Bateman [1], we state that a tree $\mathscr{T}\subset\mathscr{B}$ is lacunary of order 0 if $\mathscr{T}$ consists of a single ray containing 0, and that $\mathscr{T}$ is lacunary of order $N$ if all of the splitting vertices of $\mathscr{T}$ lie on a lacunary tree of order $N-1$. It is understood that when we say that a tree is lacunary of order $N$ (or, more colloquially, the tree is $N$-lacunary) that the tree is not lacunary of an order lower than $N$.
If $\mathscr{P}\subset\mathscr{B}$ is lacunary of order $N$ and of finite height $h(\mathscr{P})$, we say that $\mathscr{P}$ is pruned provided every ray in $\mathfrak{R}_{\mathscr{P}}(0)$ contains exactly one vertex $v_{j}$ that splits in $\mathscr{P}$ such that $\textup{split}_{\mathscr{P}}(v_{j})=j$ for $1\leq j\leq N$.
Given $\Omega\subset\mathbb{S}^{1}$, note the lacunary value $\lambda(\Omega)$ of $\Omega$ satisfies the equality
$\lambda(\Omega)=\sup\left\{N:\mathscr{T}_{\Omega}\;\textup{contains a lacunarytree of order}\;N\right\}.$
Let $0<\eta$. We say that a tree $\mathscr{T}\subset\mathscr{B}$ is $\eta$-separated provided for any splitting vertex $0a_{1}a_{2}\ldots a_{k}$ any two descendants $u$ and $v$ that are splitting vertices and lying on separate halves of the interval $Q_{0a_{1}a_{2}\ldots a_{k}}$ must be such that the Euclidean distance between the intervals $Q_{u}$ and $Q_{v}$ is greater than or equal to $\eta$ times the length of the interval $Q_{0a_{1}a_{2}\ldots a_{k}}$.
We are now in position to state the main theorem of the paper.
Theorem 1.
Let $\Omega\subset\mathbb{S}^{1}$. Suppose there exists $\eta>0$ so that, for every natural number $N$, the tree $\mathscr{T}_{\Omega}$ contains an $\eta$-separated subtree that is lacunary of order $N$. Then the maximal operator $M_{\Omega}$ is unbounded on $L^{p}(\mathbb{R}^{2})$ for every $1\leq p<\infty$.
For any natural number $N$, a function $f:\mathscr{B}^{N}\rightarrow\mathscr{B}^{N}$ is called a sticky map provided $h(f(u))=h(u)$ for every $u\in\mathscr{B}^{N}$ and moreover such that $f(u)$ is an ancestor of $f(v)$ whenever $u$ is an ancestor of $v$.
Let $\mathscr{T}\subset\mathscr{B}$ be a tree of finite height whose vertices consist of a collection of vertices $\{v_{j}\}$, all of height $h(\mathscr{T})$, together with all of the ancestors of these vertices. To every sticky map $\sigma:\mathscr{B}^{h(\mathscr{T})}\rightarrow\mathscr{T}$ we may associate a set $K_{\sigma}\subset\mathbb{R}^{2}$ defined as follows.
Let $d_{\sigma,0j_{1}\ldots j_{h(\mathscr{T})}}\in[0,1]$ denote the left-hand endpoint of the interval $Q_{0k_{1}\ldots k_{h(\mathscr{T})}}$, where
$\sigma(0j_{1}j_{2}\ldots j_{h(\mathscr{T})})=0k_{1}k_{2}\ldots k_{h(\mathscr{T})}\;.$
We let $\rho_{\sigma,0j_{1}\ldots j_{h(\mathscr{T})}}$ denote the interior of the union of all lines in $\mathbb{R}^{2}$ passing through the interval $\{0\}\times Q_{0j_{1}\ldots j_{h(\mathscr{T})}}$ oriented in the direction $(1,d_{\sigma,0j_{1}\ldots j_{h(\mathscr{T})}})\;.$ We define
$K_{\sigma}=\bigcup_{j_{1},\ldots,j_{h(\mathscr{T})}\atop j_{i}\in\{0,1\}}\rho_{\sigma,0j_{1}\ldots j_{h(\mathscr{T})}}\;.$
We set
$K_{\sigma,1}=K_{\sigma}\cap\left\{(x_{1},x_{2})\in\mathbb{R}^{2}:0\leq x_{1}\leq 1\right\}$
and
$K_{\sigma,2,\eta}=K_{\sigma}\cap\left\{(x_{1},x_{2})\in\mathbb{R}^{2}:\frac{1}{2\eta}\leq x_{1}\leq\frac{1}{2\eta}+1\right\}\;.$
Table 1 and Figure 1 provide an example of a sticky map $\sigma:\mathscr{B}^{4}\rightarrow\mathscr{B}^{4}$ and the associated set $K_{\sigma}$.
Table 1. Sticky map $\sigma:\mathscr{B}^{4}\rightarrow\mathscr{B}^{4}$
$t$ $\sigma(t)$ $t$ $\sigma(t)$
$0$ 1100 1000 $0$
0001 1101 1001 0001
0010 1110 1010 0011
0011 $1$ 1011 0010
0100 1011 1100 0100
0101 1010 1101 0100
0110 1010 1110 0101
0111 1010 $1$ 0100
Figure 1. $K_{\sigma}$ set associated to Table 1
Lemma 1 (Bateman [1]).
Suppose $\mathscr{P}\subset\mathscr{B}$ is a pruned tree that is lacunary of order $N$ and of finite height $h(\mathscr{P})$. Moreover suppose $\mathscr{P}$ contains $2^{N}$ vertices of height $h(\mathscr{P})$. Then
$|K_{\sigma,1}|\gtrsim\frac{\log N}{N}\;$
holds for every sticky map $\sigma:\mathscr{B}^{h(\mathscr{P})}\rightarrow\mathscr{P}$.
Lemma 2.
Suppose $\mathscr{P}\subset\mathscr{B}$ is an $\eta$-separated pruned tree that is lacunary of order $N$ and of finite height $h(\mathscr{P})$. Moreover suppose $\mathscr{P}$ contains $2^{N}$ vertices of height $h(\mathscr{P})$. Then there exists a sticky map $\sigma:\mathscr{B}^{h(\mathscr{P})}\rightarrow\mathscr{P}$ such that
$|K_{\sigma,2,\eta}|\lesssim_{\eta}\frac{1}{N}\;.$
Note Lemma 1 is essentially Claim 7A of [1]. Moreover, Lemma 2, without the separation assumption, corresponds to Claim 7(B) of [1]. To prove Theorem 1 it suffices to prove Lemma 2. To see this, suppose the hypotheses of Theorem 1 are satisfied. By Bateman’s pruning argument in Section 3 of [1], we have that, given $N>0$, there exists a pruned tree $\mathscr{P}\subset\mathscr{T}_{\Omega}$ of finite height $h(\mathscr{P})$ with $2^{N}$ vertices of height $h(\mathscr{P})$ that is lacunary of order $N$. As $\eta$-separation is preserved under pruning, we have $\mathscr{P}$ is $\eta$-separated. Lemma 2 provides the existence of a sticky map $\sigma:\mathscr{B}^{h(\mathscr{P})}\rightarrow\mathscr{P}$ such that the associated sets $K_{\sigma,1}$ and $K_{\sigma,2,\eta}$ satisfy $|K_{\sigma,1}|\gtrsim\frac{\log N}{N}\;$ and $|K_{\sigma,2,\eta}|\lesssim_{\eta}\frac{1}{N}\;.$ Note that the average of $\chi_{K_{\sigma,2,\eta}}$ over any parallelogram $\rho_{\sigma 0j_{1}\ldots j_{h}(\mathscr{T})}\cap\{(x,y)\in\mathbb{R}^{2}:0\leq x\leq\frac{1}{2\eta}+1\}$ equals $\frac{1}{(\frac{1}{2\eta}+1)}$. The slope $d_{\sigma 0j_{1}\ldots j_{h}(\mathscr{T})}$ of this parallelogram of width $2^{-h(\mathscr{T})}$ is within $2^{-h(\mathscr{T})}$ of a direction in $\Omega$, and accordingly is contained in a parallelogram of $\frac{1}{(\frac{1}{2\eta}+1)}$ times its area but oriented in a direction in $\Omega$. Hence $M_{\Omega}\chi_{K_{\sigma,2,\eta}}\gtrsim\eta^{2}$ on $K_{\sigma,1}$, and we have the proof of Theorem 1.
3. Proof of Lemma 2
Proof of Lemma 2.
We may assume without loss of generality that $\eta=2^{-j}$ for some natural number $j$.
Let $(x,y)\in\mathbb{R}^{2}$ with $\frac{1}{2\eta}<x\leq\frac{1}{2\eta}+1$. By linearity of expectation, it suffices to show that $Pr((x,y)\in K_{\sigma})\lesssim_{\eta}\frac{1}{N}$, where the probability is taken over all sticky maps $\sigma:\mathscr{B}^{h(\mathscr{P})}\rightarrow\mathscr{P}$.
Let $q_{1},\ldots,q_{k},\ldots,q_{2^{N}}$ denote the $2^{N}$ vertices in $\mathscr{P}$ of height $h(\mathscr{P})$. For each $1\leq k\leq 2^{N}$, we let $0q_{k1}\ldots q_{kh(\mathscr{P})}$ denote the binary string of $q_{k}$. Let $b_{1},\ldots,b_{l}$ denote all of the vertices in $\mathscr{B}$ of height $h(\mathscr{P})$ such that, if $b_{k}$ is the string $0{b_{k}}_{1}\ldots{b_{k}}_{h(\mathscr{P})}$, then for some $q_{n}$ there exists a parallelogram $\rho_{k}$ that contains $(x,y)$ with longest sides of slope $\sum_{j=1}^{h(\mathscr{P})}2^{-j}{q_{n}}_{j}$ and with corners at $(0,\sum_{j=1}^{h(\mathscr{P})}2^{-j}{b_{k}}_{j})$ and $(0,\sum_{j=1}^{h(\mathscr{P})}2^{-j}{b_{k}}_{j}+2^{-h(\mathscr{P})})$ and with a right vertical side on the line $x=\frac{1}{2\eta}+1$. Note $0\leq l=l(\mathscr{P},x,y)\leq 2^{N}$. We assume without loss of generality that $(x,y)$ does not lie on the boundary of this parallelogram and hence there is at most one parallelogram $\rho_{k}$ satisfying this property. (The set of points lying on the boundaries of all parallelograms of this form is of measure 0.)
Let $g_{0}$ denote the splitting vertex of $\mathscr{P}$ of lowest height. Let $g_{00}$ denote the splitting vertex of $\mathscr{P}$ of lowest height that is or is a descendant of one child of $g_{0}$ and we let $g_{01}$ denote the splitting vertex of $\mathscr{P}$ of lowest height that is or is a descendant of the other child. (The reader should keep in mind that, although $g_{0}$ is splitting, it is possible that one or neither of its immediate children are splitting, and for this argument we need to insure that $g_{00}$ and $g_{01}$ lie on “opposite sides of the family tree” rooted at $g_{0}$.) More generally suppose $g_{0a_{1}\ldots a_{j}}$ has been defined for $j\leq N-2$. We let $g_{0a_{1}\ldots a_{j}0}$ denote the splitting vertex of $\mathscr{P}$ of lowest height that is or is a descendant of one child of $g_{0a_{1}\ldots a_{j}}$ and we let $g_{0a_{1}\ldots a_{j}1}$ denote the splitting vertex of $\mathscr{P}$ of lowest height that is or is a descendant of the second child. Note the heights of both $g_{0a_{1}\ldots a_{j}0}$ and $g_{0a_{1}\ldots a_{j}1}$ are greater than the height of $g_{0a_{1}\ldots a_{j}}$ and they do not have to be equal to each other.
We now consider a splitting vertex $g_{0r_{1}\ldots r_{k}}$ of $\mathscr{P}$. The set of real numbers $t$ such that there exits a line passing through $(x,y)$ and $(0,t)$ with slope lying in the interval $g_{0r_{1}\ldots r_{k}}$ forms an interval $I_{g_{0r_{1}\ldots r_{k}}}$ of length less than $\frac{2}{\eta}$ times the length of the interval $g_{0r_{1}\ldots r_{k}}$. This interval is in turn contained in a union of at most $\frac{2}{\eta}+1$ dyadic intervals of length that of $g_{0r_{1}\ldots r_{k}}$ all of which intersecting $I_{g_{0r_{1}\ldots r_{k}}}$. We label those intervals in [0,1] that happen to contain any of the intervals $b_{1},\ldots,b_{l}$ as $G^{m}_{g_{0r_{1}\ldots r_{k}}}$ where $m$ is an index over a set that is possibly empty but could have integer values from $1$ to as large as $\frac{2}{\eta}+1\leq\frac{4}{\eta}$. Since $\mathscr{P}$ satisfies an $\eta$-separability condition, if $0s_{1}\ldots s_{j}\neq 0t_{1}\ldots t_{j}$ no interval $G^{m}_{g_{0s_{1}\ldots s_{j}}}$ for $11\leq j\leq N-$ 1 can be an interval $G^{n}_{g_{0t_{1}\ldots t_{j}}}$.
Note that if $g_{0a_{1}\ldots a_{j}a_{j+1}}$ is a descendant of $g_{0a_{1}\ldots a_{j}}$, then any interval $G^{m}_{g_{0a_{1}\ldots a_{j}a_{j+1}}}$ must be contained in an interval $G^{n}_{g_{0a_{1}\ldots a_{j}}}$ for some $n$. Moreover, the interval $G^{k}_{g_{0a_{1}\ldots a_{j}}}$ can contain at most $\frac{4}{\eta}$ intervals $G^{m}_{g_{0a_{1}\ldots a_{j}0}}$ and at most $\frac{4}{\eta}$ intervals $G^{m}_{g_{0a_{1}\ldots a_{j}1}}$. At this stage we recall that an $n$-ary tree is that same as a binary tree except that each vertex may have up to $n$ descendants. For example, 3-ary (or ternary) trees are considered in [2].
We define an $\frac{8}{\eta}$-tree $\mathscr{G}$ described as follows. We assume without loss of generality that $G^{k}_{g_{0}}$ exists for at least one value of $k$, as otherwise $(x,y)\notin K_{\sigma}$ for every sticky map $\sigma:\mathscr{B}^{h(\mathscr{P})}\rightarrow\mathscr{P}$ automatically holds. The root of $\mathscr{G}$ is the interval $[0,1]$. Vertices of $\mathscr{G}$ of height $h$ are the intervals $G^{k}_{g_{0a_{1}\ldots a_{h}}}$, and edges are placed between $[0,1]$ and the intervals $G^{k}_{g_{0}}$ and also between any $G^{m}_{g_{0a_{1}\ldots a_{j}}}$ and any $G^{n}_{g_{0a_{1}\ldots a_{j}a_{j+1}}}$. $\mathscr{G}$ has height $N-1$.
Note that the number of vertices in $\mathscr{G}$ of height $j$ is bounded by $\frac{4}{\eta}2^{j}$.
The $\eta$-separation condition on $\mathscr{P}$ manifests itself at this stage of the argument in a very important way. Namely, there do not exist two intervals $G^{m}_{g_{0a_{1}\ldots a_{j}0}}$, $G^{n}_{g_{0a_{1}\ldots a_{j}1}}$ that lie in the same half of an interval $G^{k}_{g_{0a_{1}\ldots a_{j}}}$. The reason for this is that, letting $H$ denote a half of the interval $G^{k}_{g_{0a_{1}\ldots a_{j}}}$, the sets
$\left\{(u,v)\in\mathbb{R}^{2}:\frac{1}{2\eta}<u,v=mu+b,b\in H,m\in g_{0a_{1}\ldots a_{j}0}\right\}$
and
$\left\{(u,v)\in\mathbb{R}^{2}:\frac{1}{2\eta}<u,v=mu+b,b\in H,m\in g_{0a_{1}\ldots a_{j}1}\right\}$
are disjoint, and hence both cannot simulaneously contain the point $(x,y)$. Figures 2 and 3 illustrate the associated disjointness.
$I$$\eta|I|$
Figure 2. $\eta$-separation condition
Figure 3. Disjointness past $x=\frac{1}{2\eta}$ associated to $\eta$-separation
Accordingly, if $\sigma:\mathscr{B}^{h(\mathscr{P})}\rightarrow\mathscr{P}$ is a randomly chosen sticky map, the probability that $(x,y)\in K_{\sigma}$ is bounded by the probability of a Bernoulli $(\frac{1}{2})$ percolation on $\mathscr{G}$, which is a subtree of an $\frac{8}{\eta}$-tree of height $N-1$ with at most $\frac{4}{\eta}2^{j}$ vertices of height $j$. Indeed, for $(x,y)$ to lie in the associated $K_{\sigma}$, a “right choice” for $\sigma$ must be made for a sequence of nested half-intervals associated to the vertices in a ray in $\mathscr{G}$, a right choice on a half-interval associated to a vertex in $\mathscr{G}$ being that, once made, there exists a sticky map $\psi:\mathscr{B}^{h(\mathscr{P})}\rightarrow\mathscr{P}$ agreeing with $\sigma$ on that vertex such that $(x,y)\in K_{\psi}$ and $K_{\psi}$ intersects the $y$-axis on an interval that is a subset of the interval associated to that vertex. Note two choices for $\sigma$ are possible for every half-interval associated to a vertex in a ray in $\mathscr{G}$, but by the disjointness properties associated to the sets displayed in the previous paragraph at most one is a right one. Note that, without the separation condition, the above disjointness property does not hold and there could be more than one right choice. As calculations similar to those found in [2, 8] indicate (see also Exercise 5.52 (a) of [9]) this probability of this Bernoulli percolation is bounded by $\frac{C}{\eta}\frac{1}{N}$, and so the lemma holds. ∎
4. An example
Fix a natural number $N$ and $(x,y)\in\mathbb{R}^{2}$. Let $\mathscr{T}$ be a pruned tree of bounded height $h(\mathscr{T})$ with lacunary order $N$. We set $\textup{Pr}_{\mathscr{T}}(x,y)$ to be the probability over all sticky maps $\sigma:\mathscr{B}^{h(\mathscr{T})}\rightarrow\mathscr{T}$ that $(x,y)\in K_{\sigma}$. In Claim 7(B) of [1] it is asserted that $\textup{Pr}_{\mathscr{T}}(x,y)\lesssim\frac{1}{N}$ provided $(x,y)\in[1,2]\times[0,3]$, although this is not necessarily the case. An example is provided by the following theorem.
Theorem 2.
Given a natural number $N$, there exists a pruned tree $\mathscr{P}$ of lacunary order $N$ and a point $(1,y)$ with $21\leq y\leq 3/$ 2 such that $Pr_{\mathscr{P}}(1,y)=1$.
Proof.
We define the interval maps $\rho_{1}$ and $\rho_{2}$ on the set of closed intervals in $\mathbb{R}$ of positive finite measure by the following:
$\rho_{1}([a,b])=\left[\frac{a+b}{2}-\frac{b-a}{8},\frac{a+b}{2}\right]\textup{;}$
$\rho_{2}([a,b])=\left[\frac{a+b}{2},\frac{a+b}{2}+\frac{b-a}{8}\right]\;.$
Let $\mathcal{I}_{0}=[0,1]$. We set $\mathcal{I}_{01}=\rho_{1}\mathcal{I}_{0}$ and $\mathcal{I}_{02}=\rho_{2}\mathcal{I}_{0}$. Similarly, for any sequence $a_{1},a_{2},\ldots,a_{k}$ of 1’s and 2’s, we let
$\mathcal{I}_{0a_{1}a_{2}\ldots a_{k}}=\rho_{a_{k}}\rho_{a_{k-1}}\ldots\rho_{a_{1}}\mathcal{I}_{0}\;.$
We let $m_{0a_{1}a_{2}\ldots a_{k}}\in[0,1]$ denote the left hand endpoint of $\mathcal{I}_{0a_{1}a_{2}\ldots a_{k}}$ Given $N$, let $\Omega_{N}$ be the set of all points of the form $m_{0a_{1}a_{2}\ldots a_{N}}+2^{-100N}$. Let $\mathscr{P}$ be the pruned tree of lacunary order $N$ and of height $3N$ consisting of all vertices in $\mathscr{B}$ corresponding to the intervals $\mathcal{I}_{0a_{1}a_{2}\ldots a_{k}}$ for $0\leq k\leq N$. Note $\mathscr{P}=\mathscr{T}_{\Omega_{N}}^{3N},$ the tree $\mathscr{T}_{\Omega_{N}}$ truncated at height $3N$.
Let now $y=1+\frac{1}{8}+(\frac{1}{8})^{2}+\cdots+(\frac{1}{8})^{N}-(\frac{1}{8})^{N+1}$. We will show that for every sticky map $\sigma:\mathscr{B}^{3N}\rightarrow{\mathscr{P}}$ we have $(1,y)\in K_{\sigma}$, proving the desired result. In particular, we will see that the probability of $(1,y)$ lying in $K_{\sigma}$ does not correspond to the survivor probability of a Bernoulli $(\frac{1}{2})$ percolation of a binary tree of height $N$, as rather for this particular value of $y$ each sticky map $\sigma$ provides a “tournament” for which exactly one slope in $\Omega_{N}$ is associated to a parallelogram in $K_{\sigma}$ that contains $(1,y)$. This is because, for every natural number $j$ and sequence $a_{1}a_{2}\ldots a_{j}$, the translates of both $\mathcal{I}_{0a_{1}a_{2}\ldots a_{j}1}$ and $\mathcal{I}_{0a_{1}a_{2}\ldots a_{j}2}$ by $\frac{1}{8}+\dots+(\frac{1}{8})^{j+1}$ lie in the same half of a dyadic interval in $[0,1]$ whose length is the same as that of the interval $\mathcal{I}_{0a_{1}a_{2}\ldots a_{j}}$.
We now discuss the above remarks in detail. It will be helpful to associate to any set $\mathcal{S}$ in $[0,1]$ its reflection across $\frac{1}{2}$ that we denote by $rf(\mathcal{S})$; in particular
$rf(\mathcal{S})=\{1-x:x\in\mathcal{S}\}\;.$
We will associate to each $\mathcal{I}_{0a_{1}a_{2}\ldots a_{k}}$ its reflection $\mathcal{J}_{0a_{1}a_{2}\ldots a_{k}}$ given by
$\mathcal{J}_{0a_{1}a_{2}\ldots a_{k}}=rf(\mathcal{I}_{0a_{1}a_{2}\ldots a_{k}})\;.$
If $h\in\mathbb{R}$ and $\mathcal{S}\subset\mathbb{R}$, we define the translate $\tau_{h}\mathcal{S}$ by $\chi_{\tau_{h}\mathcal{S}}(x)=\chi_{\mathcal{S}}(x-h)$.
[…]
$\mathcal{K}_{0a_{1}\ldots a_{N}}=\tau_{-1+y+(\frac{1}{8})^{N+1}}\mathcal{J}_{0a_{1}\ldots a_{N}}\;.$
We also define for any string $a_{1}\ldots a_{j}$ of $1$’s and $2$’s for $1\leq j\leq N$ the set
$\mathcal{K}_{0a_{1}a_{2}\ldots a_{j}}=\tau_{\sum_{k=1}^{j}2^{-3k}}\mathcal{J}_{0a_{1}a_{2}\ldots a_{j}}\;.$
Note that for $11\leq j\leq N-$ 1 both $\mathcal{K}_{0a_{1}a_{2}\ldots a_{j}1}$ and $\mathcal{K}_{0a_{1}a_{2}\ldots a_{j}2}$ lie on the right half of $\mathcal{K}_{0a_{1}a_{2}\ldots a_{j}}$
Our choice of $\Omega_{N}$ and $y$ gives the possible set $\textup{Poss}(1,y)$ a particular structure that we now wish to exploit.
Let us now fix a sticky map $\sigma:\mathscr{B}^{3N}\rightarrow{\mathscr{P}}$ We need to show that $(1,y)$ must lie in $K_{\sigma}$. Readers familiar with Bateman’s terminology might at this point observe that since $[0,1]$ and all intervals of the form $\mathcal{I}_{0a_{1}a_{2}\ldots a_{j}}$ for $11\leq j\leq N-$ 1 correspond to splitting vertices of $\mathscr{P}$, we have that $[0,1]$ and intervals of the form $\mathcal{K}_{0a_{1}a_{2}\ldots a_{j}}$ for $11\leq j\leq N-$ 1 are associated to choosing vertices of $\langle(1,y)\rangle$.
Associated to $\sigma$ will be the string of 1’s and 2’s that are defined recursively as follows.
Let $k_{1}$ be such that $\sigma[\frac{1}{2},1]\supset\mathcal{I}_{0k_{1}}\;.$ Note that $\sigma(\mathcal{K}_{0k_{1}})=\mathcal{I}_{0k_{1}}$ because ${\mathscr{P}}$ has no splitting vertices between $\mathcal{I}_{0}$ and $\mathcal{I}_{0k_{1}}$.
Assuming $k_{1},\ldots,k_{j}$ are determined, let $k_{j+1}$ be the value such that
$\sigma(\textup{right half }(\mathcal{K}_{0k_{1}k_{2}\ldots k_{j}}))\;\supset\;\mathcal{I}_{0k_{1}k_{2}\ldots k_{j+1}}\;.$
Note that $\sigma(\mathcal{K}_{0k_{1}k_{2}\ldots k_{j+1}})=\mathcal{I}_{0k_{1}k_{2}\ldots k_{j+1}}$ since there are no splitting vertices in ${\mathscr{P}}$ between $\mathcal{I}_{0k_{1}k_{2}\ldots k_{j}}$ and $\mathcal{I}_{0k_{1}k_{2}\ldots k_{j+1}}$.
We have that $(1,y)$ lies in the parallelogram in $K_{\sigma}$ with slope $m_{0k_{1}\ldots k_{N}}$ whose left hand side is the vertically oriented interval $\{(0,t):t\in\mathcal{K}_{0k_{1}\ldots k_{N}}\}$. ∎
5. Concluding remarks
Even though Theorem 2 provides a counterexample to a step in the proof of Theorem 1 of [1], it does not provide a counterexample to Theorem 1 of [1] itself as the union of all of the trees $\mathscr{T}_{\Omega_{j}}$ in the proof of Theorem 2 contains, for every natural number $N$, a subtree that is $\frac{1}{8}$-separated and lacunary of order $N$.
Theorem 1 of [1] was used as a key step in the proofs of Theorems 2,3, and 4 of [5] and Theorem 4 of [6]. Claim 7(B) of [1] was used in the proofs of Theorems 2.1 and 2.2 of [4] and Theorem 1 of [7] . Researchers in the area should consider these results at the moment to be at best provisional.
It is important to recognize that a set $\Omega\subset[0,1]$ may have infinite lacunary value yet be such that for no $\eta>0$ does $\mathscr{T}_{\Omega}$ contain an $N$-lacunary $\eta$-separated subtree for every $N$. Such a set may be defined as follows.
Let $j\in\mathbb{N}$. We define the interval maps $\rho_{j,1},\rho_{j,2}$ respectively on the set of closed intervals in $\mathbb{R}$ of finite length by
$\rho_{j,1}([a,b])=\left[\frac{a+b}{2}-2^{-j}(b-a),\frac{a+b}{2}\right]\;,$
$\rho_{j,2}([a,b])=\left[\frac{a+b}{2},\frac{a+b}{2}+2^{-j}(b-a)\right]\;.$
Let $\Omega\subset[0,1]$ be the set of all $x$ contained in infinitely many intervals of the form
$\rho_{1,i_{1}}\rho_{2,i_{2}}\cdots\rho_{k,i_{k}}[0,1]\;.$
One can check that $\lambda(\Omega)=\infty$. However, there is no $\eta>0$ for which $\mathscr{T}_{\Omega}$ contains an $N$-lacunary $\eta$-separated subtree for every $N$.
This set $\Omega$ constructed here poses a model problem for the theory of geometric maximal operators. Is the geometric maximal operator $M_{\Omega}$ bounded on $L^{p}(\mathbb{R}^{2})$ for every $1<p\leq\infty$